Present: Lothar, Ian, Michael, Peter, Alessandra, StefanoL, Tony, Nick Excused: StefanoB

Agenda: - action items - PhEDEx development plans and priorities (mail from Lassi) - if/how/when to engage on proving Data Management for "detectors", i.e. test beam, MTCC, cosmic tests, tracker test.. should they use DBS, PhEDEx ? (e.g. tracker just announced they will likely write up to 100TB by March 07) - integration checklist - subproject status reports - AOB

Action Items

  • [Ian] to talk to Lucia
    • Ian: CMSSW 0_7_2 should be ok for the val samples

  • [Ale] DBS contaisn DS needed by PRS parts not in dbs or badly imported
    • Luciano: information present on zipdb but it doesn't include 10% of DS in Darin's list Problem reasonably understood. Luciano to sort it out with Lassi.

  • [Pete] local site conf with Ian Michael. Ian needs the SE name syntax agreed on. CLOSED

PhEDEx development plans and priorities (mail from Lassi)

  • 8th to jul 31st Lassi is away

  • [pX] are as referenced in Lassi's email

  • [p1] Jens, Dimitrije around phedex. Dimtrije will be away beginning of August. Operations together with Jens and Daniele. Michael at cern 17/7 - 24/7. Daniele at cern up to 12/13 th of July. Daniele to take over from Lassi, get next to Lassi after his return.

  • [p2] Michael as point of contact for p2

  • [p4] and [p5] for the time being manual injection to dbs implied operations model depends on xfers to cern, doing the bookkeeping development (to start after Lassi is back)

  • Pete Lassi to touch base tomorrow on the phedex plans

if/how/when...

... to engage on proving Data Management for "detectors", i.e. test beam, MTCC, cosmic tests, tracker test.. should they use DBS, PhEDEx ? (e.g. tracker just announced they will likely write up to 100TB by March 07)

  • Not discussed

integration checklist

  • what are the next steps for the 50Mevts production? Ian: reasonable shape to exercise cpu not to run the 50m. 15 July is the release date for the 080 version need to find a person to follow up the out of T2 subscriptions and file cleanup

  • op teams: 2 teams in place, a 3rd to rump up. New people to join in with training provided by running val samples

  • Ian: hitting bottleneck of RB -> will not hit the 50k/jobs per day with the current infrastructure for csa '06

subproject status reports

  • Lee: dbs participation in wm workshop. Issues with merge step. Preparation for analysis workshop (agenda?)
  • Tony: input buffer tests eb emulator needs Lassi's time for updating phedex dropbox on lxgate10. Done

-- NickJSinanis - 11 Jul 2006

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-07-11 - NickJSinanis
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback